OpenStudy (anonymous):

Solve the equation -3y - 8=7y + 22

OpenStudy (anonymous):

Isolate y on one side first. Go from here: 10y+22=-8

OpenStudy (anonymous):

i just add 10y+22 right?

OpenStudy (anonymous):

You should subtract 22 from both sides first so you'd have 10y=__, then all you have to do is divide by 10.

OpenStudy (anonymous):

-3y-8=7y+22 -8=10y+22 -8-22=10y

OpenStudy (anonymous):

divided by 10y?

OpenStudy (anonymous):

Once you get to -30=10y, all you have to do is divide both sides by 10. This will put y by itself and you'll be done.

OpenStudy (anonymous):

and i get -3

OpenStudy (anonymous):

Yep. That's it.

OpenStudy (anonymous):

y=-3

OpenStudy (anonymous):

thank you